OpenStudy (anonymous):

Complete the table of values for the following equation. 8x + 6y = 26 x-values y-values -5 11 7 1 3 4 7 -5 10 Select the best answer from the choices provided. x-values y-values -5 11 -2 7 1 3 4 -1 7 -5 10 -9 x-values y-values -5 11 -3 7 1 3 4 -2 7 -5 10 -9 x-values y-values -5 11 -2 7 1 3 4 -2 7 -5 10 -7 x-values y-values -5 11 -1 7 1 3 4 -3 7 -5 10 -7

OpenStudy (texaschic101):

just a matter of subbing 8x + 6y = 26.....when x = 7 8x + 6y = 26...when x = 4 8(7) + 6y = 26 8(4) + 6y = 26 56 + 6y = 26 32 + 6y = 26 6y = 26 - 56 6y = 26 - 32 6y = - 30 6y = - 6 y = -5 y = -1 so when x = 7, y = -5 so when x = 4, y = -1 try another number.. 8x + 6y = 26.....when x = 10 8x + 6y = 26...when x = -2 8(10) + 6y = 26 8(-2) + 6y = 26 80 + 6y = 26 -16 + 6y = 26 6y = 26 - 80 6y = 26 + 16 6y = - 54 6y = 42 y = -9 y = 7 so when x = -2, y = 7 so when x = 10, y = -9 so which one has numbers that match ^^ these numbers ??